# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v1

@@ 3.2 Two-Stage PG-DPO Algorithm
3.3 Rationale for the Two-Stage Approach: Leveraging Costate Estimation
The preceding subsections introduced two variants of the PG-DPO framework: a baseline approach (§3.1) and a Two-Stage approach (§3.2). A natural question arises: why introduce the Two-Stage variant, which potentially adds complexity, instead of solely relying on optimizing the policy networks directly? This subsection elaborates on the theoretical and practical motivations.
3.3.1 Qualitative Rationale: Exploiting PMP Structure and Costate Properties
At its core, the rationale stems from the structure provided by Pontryagin’s Maximum Principle (PMP) itself and the economic meaning of the costate variables. As established in Section 2.2, PMP characterizes the optimal controls ($\bm{\pi}_{t}^{*},C_{t}^{*}$) directly in terms of the costates ($\lambda_{t}^{*},\bm{\eta}_{t}^{*}$) and their relevant derivatives (e.g., $\partial_{x}\lambda_{t}^{*},\partial_{\mathbf{Y}}\lambda_{t}^{*}$). The analytical formulas used in Stage 2 of the Two-Stage method are precisely these first-order conditions. Therefore, if accurate estimates of these costates and their derivatives could be obtained, PMP provides a direct blueprint for constructing the optimal policy.
Furthermore, the primary costate $\lambda_{t}$ represents the marginal utility of wealth ($\lambda_{t}\approx\partial V/\partial x$). Economic principles, like consumption smoothing, suggest agents strive for a stable marginal utility path. Mathematically, this often means the value function $V$ and its derivatives ($\lambda_{t},\partial_{x}\lambda_{t},\partial_{\mathbf{Y}}\lambda_{t}$, etc.) exhibit relatively smooth behavior or possess a simpler analytical structure compared to the optimal policy functions $(\bm{\pi}^{*},C^{*})$ themselves. The optimal policy might need to be highly complex, especially in non-affine settings or under constraints.
This difference in potential complexity has direct implications for learning dynamics. It is plausible that learning the potentially smoother or more structured functions associated with the costates via BPTT during the warm-up phase (Stage 1) is an easier task-requiring fewer iterations for stabilization-than learning the full complexity of the optimal policy networks $(\bm{\pi}_{\theta},C_{\phi})$ until convergence. This aligns with empirical observations that costate estimates often stabilize relatively quickly.
The potential for faster convergence or stabilization of costate-related functions compared to the full policy motivates the Two-Stage approach. This is further encouraged, albeit heuristically for general models, by theoretical stability results known for specific structured models, as formalized in the next subsection.
3.3.2 Derivative Stability in Structured Models
For certain classes of portfolio choice problems that admit explicit solutions, such as those within affine or quadratic frameworks, a strong link exists between the accuracy of the value function approximation and the accuracy of its derivatives, provided the approximation retains the original structure. This property is formally stated as follows:
Proposition 3.1 (Derivative Closeness from Value Closeness in Structured Models).
Consider a class of portfolio choice models (e.g., Merton with HARA utility (Merton, , 1971), Kim & Omberg with HARA utility (Kim & Omberg, , 1996), or Liu’s affine/quadratic framework with CRRA utility (Liu, , 2007)) where the optimal value function $V^{*}(t,w,\mathbf{y})$ has a known analytical structure (e.g., separable or exponential-polynomial in state variables with coefficients determined by ODEs). Let $\tilde{V}(t,w,\mathbf{y})$ be another function defined on a compact domain $\mathcal{D}\subset[0,T]\times\mathbb{R}^{+}\times\mathbb{R}^{k}$ that shares the same analytical structure as $V^{*}$. If the value functions are uniformly close on $\mathcal{D}$, i.e.,
$\sup_{(t,w,\mathbf{y})\in\mathcal{D}}|V^{*}(t,w,\mathbf{y})-\tilde{V}(t,w,\mathbf{y})|\longrightarrow 0,$
then their corresponding spatial derivatives (first and second order with respect to $w$ and $\mathbf{y}$) also converge uniformly on $\mathcal{D}$:
$\sup_{\mathcal{D}}\|\nabla_{w,\mathbf{y}}V^{*}-\nabla_{w,\mathbf{y}}\tilde{V}\|\to 0\quad\text{and}\quad\sup_{\mathcal{D}}\|\nabla^{2}_{w,\mathbf{y}}V^{*}-\nabla^{2}_{w,\mathbf{y}}\tilde{V}\|\to 0.$
(Proof sketch provided in Appendix C for the Liu (2007) framework).
Implication for PG-DPO Rationale.
Proposition 3.1 highlights that for well-structured problems (like those within the affine class for which the proposition holds), the derivatives inherent in the PMP conditions are well-behaved with respect to approximations that preserve the specific analytical structure. However, a key motivation for employing methods like PG-DPO is their potential applicability to a broader class of problems, including models that may lack such convenient structures — for instance, models involving non-affine dynamics (such as the CEV examples considered later in this paper). The true optimal value function $V^{*}$ for these more general models may not possess the specific separable or exponential-polynomial form assumed in the proposition.
Consequently, Proposition 3.1, proven for specific structured models, does not directly guarantee that similar derivative stability holds for the wider range of potentially non-affine models that PG-DPO aims to tackle. Establishing whether analogous stability properties exist in these more general settings, or rigorously proving the convergence of derivative estimates obtained via BPTT from neural network approximations in such cases, remains an area for future research and is beyond the scope of this paper.
Nevertheless, the proposition provides valuable context. It confirms that the target objects (optimal value function derivatives) possess inherent stability within important, albeit specific, classes of models. This offers heuristic support for the hypothesis underlying the Two-Stage PG-DPO: namely, that the costate derivatives, being fundamental quantities in a well-posed optimal control problem, might be estimated reliably via BPTT (Stage 1) and effectively used in the PMP formulas (Stage 2), even when tackling more complex, potentially non-affine models.

### v2

@@ 3.2 Two-Stage PG-DPO Algorithm
3.3 Rationale for the Two-Stage Approach
3.3.1 The Parabolic Nature of Financial Models and PDE Regularity
Table 1: Classification of Parabolic HJB PDEs in Financial Models
Class Diffusion Matrix $a(t,x,\mathbf{y})$ Property SDE Coefficient Requirements and Dynamic Portfolio Examples
(U) Uniformly Elliptic Uniformly positive definite across the state space The diffusion coefficients $\bm{\sigma}$ (for wealth $X_{t}$) and $\bm{\sigma}_{Y}$ (for state $\mathbf{Y}_{t}$) are such that the resulting diffusion matrix in the HJB is non-zero and bounded below by a positive constant everywhere ($\bm{\sigma}\Psi\bm{\sigma}^{\top}$ and $\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}$ uniformly positive definite). Drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) are also sufficiently regular (e.g., Lipschitz).
($\implies$ Value Function / Costate very smooth) Examples: Merton’s portfolio problem with constant coefficients ($\mu,\sigma>0,r$ are constants); models with state-dependent volatility that remains strictly positive; the multi-asset OU model in Section 2.3.1 (if asset volatilities $\sigma_{i}>0$, factor volatility $\sigma_{Y}>0$, and $\Psi$ is positive definite).
(H) Hypo-elliptic / Uniform on Compacts May degenerate at some points, but regularity is achieved via system structure (e.g., drift-diffusion coupling via Hörmander’s condition) or holds within a compact domain of interest. Degeneracy occurs away from boundaries. Diffusion coefficients ($\bm{\sigma}$ or $\bm{\sigma}_{Y}$) can become zero at certain points within the domain. However, interaction with non-zero drift terms or other diffusion terms maintains regularity.
($\implies$ Value Function / Costate smooth in relevant regions) Examples: The multi-asset multi-factor model in Section 2.3.2 (if some factor volatilities in $\bm{\sigma}_{Y}$ can be zero at certain states but the overall system structure, possibly involving $\kappa_{Y}$, ensures hypo-ellipticity); a Vasicek interest rate model affecting asset returns $\mu$ but not $\sigma$.
(D) Degenerate at the Boundary Degenerates ($\to 0$) as the state approaches a boundary. The diffusion coefficients $\bm{\sigma}$ or $\bm{\sigma}_{Y}$ (or the combined diffusion matrix) approach zero as a state variable (e.g., wealth, volatility, interest rate) approaches a boundary (typically 0).
($\implies$ Value Function / Costate regularity may break down near the boundary) Examples: CIR (Cox-Ingersoll-Ross) interest rate/volatility model (diffusion term like $\sigma_{Y}\sqrt{Y_{t}}$); Heston stochastic volatility model (asset diffusion term like $\sigma\sqrt{\nu_{t}}$).
In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are modeled by stochastic differential equations (SDEs), the Value Function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically solves a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming & Soner, , 2006; Yong & Zhou, , 2012). This HJB equation exhibits a parabolic form. The specific nature of its parabolicity is primarily determined by the diffusion coefficients of the underlying SDEs, namely $\bm{\sigma}$ (related to assets) and $\bm{\sigma}_{Y}$ (related to state variables). While the drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) influence the first-order terms in the HJB and are crucial for the solution’s properties, they do not dictate the type’ of parabolicity, which is mainly governed by the second-order terms (diffusion).
The behavior of the diffusion matrix $a(t,x,\mathbf{y})$ (the matrix multiplying second-order derivatives in the HJB equation) allows for classifying parabolic PDEs into three main classes (see, e.g., Evans, , 2022). This classification significantly impacts the regularity (smoothness) of the Value Function $V$ and its associated Costates (the first-order derivatives $\nabla V$) (cf. Evans, , 2022, Ch. 6-7). Costates represent the sensitivity of the Value Function, and Pontryagin’s Maximum Principle links optimal Costates to the optimal policy. The core idea of Two-Stage PG-DPO involves estimating Costates and their derivatives via BPTT and using these estimates in the PMP formulas. Therefore, the regularity of the Costates—tied to the model’s parabolicity class—directly affects the difficulty of Costate estimation and the effectiveness of the Two-Stage method.
The table 1 summarizes these classes, their properties, links to SDE coefficients, and relevant examples within the dynamic portfolio optimization context.
Models falling into classes (U) and (H) are expected to possess sufficiently regular Costates, providing a favorable setting for the estimation required by the Two-Stage PG-DPO method. The numerical experiments in this paper are conducted using models belonging to classes (U) or (H), where the Two-Stage PG-DPO effectively leverages the expected regularity of the Costates. The method is particularly effective as it generally avoids visiting degenerate boundaries during practical implementation.
3.3.2 Mathematical Foundation: Derivative Stability via PDE Comparison
A key theoretical insight, particularly demonstrable in well-structured (e.g., affine) models, comes from analyzing the stability of value function derivatives under policy perturbations using regularity theory for parabolic partial differential equations (e.g., Evans, , 2022; Krylov, , 2024). This perspective illustrates how closeness of policies (in a specific sense) can imply closeness of derivatives, providing a foundation for why the Two-Stage approach might be effective.
First, let us recall the PDE associated with a given policy. For an admissible and sufficiently smooth policy $(\pi(t,z),c(t,z))$, where $z=(x,y)$ represents the state, the corresponding value function $V(t,z)$ (representing the expected utility achieved under this *specific* policy from state $z$ at time $t$) satisfies the linear parabolic PDE:
$-\frac{\partial V}{\partial t}=\mathcal{L}_{\pi,c}[V](t,z)+f_{\pi,c}(t,z),\quad\text{subject to }V(T,z)=g(z).$ (16)
Here, $g(z)=\kappa e^{-\delta T}U(x)$ is the terminal utility, $f_{\pi,c}(t,z)=e^{-\delta t}U(c(t,z))$ is the instantaneous utility source term under policy $c$, and $\mathcal{L}_{\pi,c}$ is the second-order linear differential generator associated with the state dynamics under policy $(\pi,c)$:
$\mathcal{L}_{\pi,c}[V]=$ $V_{x}\left[x\left((1-\pi^{\top}\mathbf{1})r(t,y)+\pi^{\top}\mu(t,y)\right)-c(t,z)\right]+V_{\mathbf{y}}^{\top}\mu_{Y}(t,y)$
$+\frac{1}{2}V_{xx}x^{2}\pi^{\top}\sigma(t,y)\Psi\sigma(t,y)^{\top}\pi$
$+V_{x\mathbf{y}}^{\top}x\pi^{\top}\sigma(t,y)\rho\sigma_{Y}(t,y)^{\top}$ (17)
$+\frac{1}{2}\text{Tr}\left(V_{\mathbf{y}\mathbf{y}}\sigma_{Y}(t,y)\Phi\sigma_{Y}(t,y)^{\top}\right)-\delta V$
(We incorporate the time discount term $-\delta V$ into the operator $\mathcal{L}_{\pi,c}$ for convenience). Let $V^{*}(t,z)$ denote the true optimal value function of the control problem defined in Section 2.1, achieved by the optimal policy $(\pi^{*}(t,z),c^{*}(t,z))$. Assuming sufficient regularity, $V^{*}$ satisfies the non-linear Hamilton-Jacobi–Bellman (HJB) equation (Fleming & Soner, , 2006), which implies it also satisfies the linear PDE (16) when the optimal policy $(\pi^{*},c^{*})$ is used:
$-\frac{\partial V^{*}}{\partial t}=\mathcal{L}_{\pi^{*},c^{*}}[V^{*}](t,z)+f_{\pi^{*},c^{*}}(t,z),\quad\text{subject to }V^{*}(T,z)=g(z).$ (18)
To formalize the notion of policy closeness relevant to the PDE coefficients, we introduce the following definition. Let $\Sigma_{\text{asset}}(t,y)=\sigma(t,y)\Psi\sigma(t,y)^{\top}$ and $\Sigma_{\text{asset-state}}(t,y)=\sigma(t,y)\rho\sigma_{Y}(t,y)^{\top}$.
Definition 1 (Coefficient-relevant $\epsilon$-closeness).
Let the spatial domain be
$\Omega\subseteq(0,\infty)\times\mathbb{R}^{k},\qquad z=(x,\mathbf{y})\in\Omega,$
i.e. $x$ is wealth and $\mathbf{y}\in\mathbb{R}^{k}$ is the $k$-dimensional state vector. (In practice $\Omega$ can be either the full space or a bounded truncated box that contains the region the process visits with high probability.)
Fix exponents $p>d+2$ and $q\geq 2$, where $d=\dim\Omega=k+1$. For any measurable function $h(t,z)$ set
[…]
Let $(\pi^{*},c^{*})$ be the optimal policy and $(\pi,c)$ any admissible policy. We say $(\pi,c)$ is $(\epsilon,p,q)$-close to $(\pi^{*},c^{*})$ on $Q_{T}=(0,T)\times\Omega$ if
$\|c-c^{*}\|_{L^{q,p}}$ $\leq\epsilon,$
$\|\pi^{\top}\mathbf{1}-\pi^{*\top}\mathbf{1}\|_{L^{q,p}}$ $\leq\epsilon,$
$\|\pi^{\top}\bm{\mu}-\pi^{*\top}\bm{\mu}\|_{L^{q,p}}$ $\leq\epsilon,$
$\|\pi^{\top}\Sigma_{\text{asset}}\pi-\pi^{*\top}\Sigma_{\text{asset}}\pi^{*}\|_{L^{q,p}}$ $\leq\epsilon,$
$\|\pi^{\top}\Sigma_{\text{asset-state}}-\pi^{*\top}\Sigma_{\text{asset-state}}\|_{L^{q,p}}$ $\leq\epsilon.$
Proposition 3.1 (Derivative closeness under coefficient $\epsilon$-closeness).
Let $V^{*}$ be the optimal value function corresponding to the optimal policy $(\pi^{*},c^{*})$ (cf. (18)) and let $V$ be the value function generated by another admissible policy $(\pi,c)$ (cf. (16)). Set
$D\;:=\;V^{*}-V,\qquad Q_{T}:=(0,T)\times\Omega.$
Assume
  1. (A1)
    Uniform parabolicity. The operator $\mathcal{L}_{\pi^{*},c^{*}}$ is uniformly parabolic on $Q_{T}$.
  2. (A2)
    Baseline regularity. There exist exponents $p>d+2$ and $q\geq 2$ (with $d=\dim\Omega=k+1$) such that the value functions lie in the Sobolev space (e.g., Adams & Fournier, , 2003; Evans, , 2022)
    $V^{*},\,V\;\in\;L^{q}\bigl{(}0,T;W^{2,p}(\Omega)\bigr{)}.$
  3. (A3)
    Coefficient $\epsilon$-closeness. The candidate policy $(\pi,c)$ is $(\epsilon,p,q)$-close to $(\pi^{*},c^{*})$ in the sense of Definition 1.
Then $D$ satisfies the linear PDE
$-\partial_{t}D\;=\;\mathcal{L}_{\pi^{*},c^{*}}[D]\;+\;S^{\prime},\quad\text{where}\quad S^{\prime}=(\mathcal{L}_{\pi^{*},c^{*}}-\mathcal{L}_{\pi,c})[V]+\bigl{(}f_{\pi^{*},c^{*}}-f_{\pi,c}\bigr{)}.$
Moreover, the source term is small in the weighted norm:
$\|S^{\prime}\|_{L^{q,p}}\;\leq\;C_{0}\,\epsilon,$
and by standard $L^{p}$ / Schauder estimates for uniformly parabolic equations (see, e.g., Evans, , 2022, Ch. 7),
$\|D\|_{L^{q}\!\bigl{(}0,T;W^{2,p}(\Omega)\bigr{)}}\;\leq\;C\,\|S^{\prime}\|_{L^{q,p}}\;\leq\;C\,\epsilon,$
for a constant C independent of $\epsilon$.
Consequently, the value-function gap and all first- and second-order spatial derivatives satisfy
$\bigl{\|}V^{*}-V\bigr{\|}_{L^{q,p}}+\bigl{\|}\nabla V^{*}-\nabla V\bigr{\|}_{L^{q,p}}+\bigl{\|}D^{2}V^{*}-D^{2}V\bigr{\|}_{L^{q,p}}\;\leq\;C\,\epsilon.$
Hence the costates $(\nabla V)$ produced under $(\pi,c)$ are $O(\epsilon)$-close to the optimal costates on the entire space–time domain.
Remark on assumptions, PG–DPO rationale, and scope of the result.
Proposition 3.1 makes precise the intuition for the uniformly–parabolic class (“U class”’): if two policies yield close PDE-coefficients in the weak sense of Definition 1 (only a global [^1] A still weaker occupation-weighted norm is discussed below. $L^{q,p}$-distance), then their value functions—and hence their gradients (costates)—differ by at most $O(\epsilon)$. Assumptions (A1)–(A2) are the classical entry ticket for $L^{p}$ / Schauder regularity (Evans, , 2022) and are easily verified for affine models or for models that become uniformly elliptic after adding a small floor to each diffusion coefficient (e.g. replace $\sigma(t,z)$ by $\sqrt{\sigma(t,z)^{2}+\varepsilon^{2}}$ with $\varepsilon\sim 10^{-4}$). Such a floor leaves the economics virtually unchanged while preventing the diffusion matrix from collapsing to zero, which is exactly what uniform parabolicity requires. The genuinely new ingredient is (A3), the $\epsilon$-coefficient closeness that Stage 1 of PG-DPO is designed to enforce.
Stage 1 produces a sub-optimal policy $(\pi,c)$ whose coefficient error is $\epsilon$. Via Proposition 3.1 this translates into an $O(\epsilon)$ error for the costate $\lambda\approx\nabla V$, so the BPTT outputs obtained during the short warm-up are already $O(\epsilon)$-accurate surrogates of the true optimal costates. Stage 2 simply substitutes these estimates into the PMP first-order conditions (Pontryagin, , 2014; Yong & Zhou, , 2012), thereby constructing a near-optimal control without having to train the policy network to full convergence.
Global vs.– local norms.– The result is stated with a global $\|\,\cdot\,\|_{L^{q,p}}$ norm for transparency. If desired, one may replace it by the occupation-weighted norm
[…]
where $\mu_{(\pi,c)}$ is the state-density under the candidate policy. Because the proof relies on local PDE estimates, the same $O(\epsilon)$ bound holds—now only on the subset of the state space that the process actually visits with non-negligible probability. This is valuable for models whose coefficients degenerate in remote regions.
Beyond the U class. – For hypo-elliptic (H class) (cf. Hörmander, , 1967) or boundary-degenerate (D class) models (e.g., Oleinik, , 2012), the full global $L^{p}$ theory is subtler, yet the algorithmic logic of Two-Stage PG-DPO is unaffected:
  • the PMP equations and BPTT apply to any differentiable dynamics;
  • costates typically stabilise faster than the full optimal control;
  • a brief warm-up that makes the coefficient error small where it matters is enough to obtain high-quality costate estimates.
Even when the rigorous guarantee of Proposition 3.1 is unavailable, this costate-first, PMP-second recipe remains an effective practical strategy—an observation confirmed empirically in the numerical section (Section 4) and worth exploring further for genuinely non-affine or degenerate models.
3.3.3 Qualitative Rationale and Broader Applicability
While the formal analysis in Proposition 3.1 provides rigorous support under specific regularity conditions (like uniform parabolicity, often associated with affine models), the core ideas underpinning the Two-Stage PG-DPO approach possess a broader qualitative appeal that suggests applicability beyond these settings.
Crucially, neither the Pontryagin Maximum Principle (PMP) itself (Pontryagin, , 2014; Yong & Zhou, , 2012) nor the mechanics of backpropagation-through-time (BPTT) (e.g., Goodfellow, , 2016) are inherently restricted to affine or uniformly parabolic systems. PMP provides necessary optimality conditions for general nonlinear dynamics, provided the required functions are differentiable. Similarly, BPTT serves as a general algorithm for computing gradient estimates through complex computational graphs derived from simulations, regardless of the underlying model’s linearity.
Furthermore, there is a compelling economic and mathematical intuition that the PMP costates ($\lambda\approx\nabla V$, representing marginal values) might exhibit properties that make them more amenable to estimation than the potentially complex optimal policy $\pi^{*}$ itself, especially in high-dimensional problems. For instance, the costate functions might be representable by smoother or lower-complexity functions compared to $\pi^{*}$. Alternatively, the process of learning or estimating these costate-related functions via BPTT with neural networks might converge more rapidly or stably than directly learning the intricate mapping defined by the optimal policy.
The Two-Stage PG-DPO strategy is designed precisely to leverage this potential advantage. It effectively decouples the optimization challenge into two parts: (i) tackling the potentially more tractable task of estimating costates (and their relevant derivatives) using BPTT during a warm-up phase (ii) employing the robust analytical structure of the PMP first-order conditions to translate these estimates into effective control actions during deployment.
Therefore, even in scenarios where the rigorous derivative closeness guarantees of Proposition 3.1 are harder to establish formally (e.g., for non-affine models, or systems falling into the hypo-elliptic or degenerate parabolic classes), this underlying qualitative rationale remains persuasive. The potential for BPTT-derived costate estimates to stabilize relatively quickly and capture essential optimality information offers a strong motivation for exploring the Two-Stage PG-DPO approach in these more complex settings.
This rationale, significantly reinforced by the numerical success demonstrated in the affine benchmark case (Section 4), positions PG-DPO (particularly the superior Two-Stage variant) as a promising and potentially powerful tool for tackling challenging high-dimensional, possibly non-affine, continuous-time control problems prevalent in finance and other domains. Naturally, its practical effectiveness in such extended settings must be carefully evaluated empirically on a case-by-case basis, highlighting an important direction for future research.